Down Under Boomerang, Inc., is considering a new 3-year expansion project that requires an initial fixed asset investment of $2.31 million. The fixed asset falls into the 3-year MACRS class (MACRS schedule). The project is estimated to generate $1,770,000 in annual sales, with costs of $668,000. The project requires an initial investment in net working capital of $370,000, and the fixed asset will have a market value of $360,000 at the end of the project.

  

a. If the tax rate is 22 percent, what is the project’s Year 0 net cash flow? Year 1? Year 2? Year 3? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and enter your answers in dollars, not millions of dollars, rounded to two decimal places, e.g., 1,234,567.89.)
b. If the required return is 12 percent, what is the project's NPV? (Do not round intermediate calculations and enter your answer in dollars, not millions of dollars, rounded to two decimal places, e.g., 1,234,567.89.)
a. Year 0 cash flow $-2,680,000.00 *answer correct*
Year 1 cash flow $1,028,943.06 *answer correct*
Year 2 cash flow $1,085,454.90 *answer correct*
Year 3 cash flow not attempted
b. NPV not attempted
Answer #1

3-year MACRS class (MACRS schedule) will have a depreciation rate of 7.41% in the fourth year. The project life is 3 years. Hence, the book value of the investment at the end of the project = Initial fixed asset investment x 7.41% = $ 2,310,000 x 7.41% = $ 171,171.00

the fixed asset will have a market value of $360,000 at the end of the project

Hence, gain = market value - book value = 360,000 - 171,171 = $ 188,829.00

Hence, tax on gain = Tax rate x Gain = 22% x 188,829.00 = $ 41,542.38

Hence, post tax salvage value = Market value - tax on gain on sale = 360,000 - 41,542.38 = $ 318,457.62

Please see the table below. Please be guided by the second column titled “Linkage” to understand the mathematics. The rows highlighted in yellow contain your answer. Figures in parenthesis, if any, mean negative values. All financials are in $. Adjacent cells in blue contain the formula in excel I have used to get the final output.

Hence, your final answers are:

Part a. the project’s Year 0 net cash flow = $ -2,680,000.00

Year 1 = $ 1,028,943.06

Year 2 = $ 1,085,454.90

Year 3 = $  1,623,282.04


b.   If the required return is 12 percent, what is the project's NPV?

NPV = $  259,437.25